# The quantum-refractive evolution of polarization states in pulsar emission

### v1

@@ 2.2.2 Approximate analytical solutions
Plotting the birefringent functions $\hat{\Omega}_{1}\left(s\right)$ and $\hat{\Omega}_{2}\left(s\right)$, as given by (18) and (19), respectively, one can observe that they feature distinctive patterns; they can be well approximated by some analytic models, whose curves resemble the original plots. In Fig. 7 are plotted the birefringent functions for the three cases: (a) $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$), (b) $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$), (c) $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$) with solid lines (see Figs. 7a, 7b and 7c, respectively), where they have been evaluated with the same initial condition as assumed in Section 2.2.1.
@@
where $a_{1,2}>0$, $b>0$ and $p>0$ are free parameters; with suitable values chosen for these, our model functions can give rise to solutions of equations (23)-(25) that match fairly well the numerical results obtained in Section 2.2.1. Here we express
$a_{1,2}=-\hat{\Omega}_{1,2\left(\mathrm{min}\right)}\left[\frac{e}{\left(1-\chi\right)s_{\dagger 1}+\chi s_{\dagger 2}}\right]^{\frac{p+1}{p}},~{}~{}b=\frac{p+1}{p\left[\left(1-\chi\right)s_{\dagger 1}+\chi s_{\dagger 2}\right]},$ (31)
where $\hat{\Omega}_{1,2\left(\mathrm{min}\right)}=\hat{\Omega}_{1,2}\left(s_{\dagger 1,2}\right)$, evaluated from (18), (19), with $s_{\dagger 1,2}$ denoting the extremum. We have set $p=100$ (a sufficiently large number) for all three cases, and $\chi=0.999$ for (a) and (b), and $\chi=0.056$ for (c) in Fig. 7.888In particular, the values for $\chi$ have been chosen such that our solutions converge to the asymptotic limits that match well the numerical results given by Figs. 3b, 4b and 5b in Section 2.2.1, as $s$ tends to $\infty$.
@@
Figure 7: Plots of the birefringent functions $\hat{\Omega}_{1}\left(s\right)$ and $\hat{\Omega}_{2}\left(s\right)$ and their approximate analytic models.
Plugging equation (30) into equations (23)-(25), we obtain analytical solutions as follows (for a complete derivation, see Appendix B):
@@
and $M_{\kappa,\mu}\left(z\right)$ denotes a Whittaker function of the first kind, [^9] The expression inside the square brackets in equation (35) has been reduced from its original form as given by equation (60) in Appendix B, using the identity $M_{\left(2p+1\right)/\left(2p\right),\left(p+1\right)/\left(2p\right)}\left(bs\right)=\left(bs\right)^{\left(2p+1\right)/\left(2p\right)}e^{-bs/2}M\left(0,\left(2p+1\right)/p,bs\right)$, with the Kummer function $M\left(0,\left(2p+1\right)/p,bs\right)=1$ as a special case (NIST (2024)). and $a_{1,2}$ and $b$ are given by (31). Here we determine $S_{\mathrm{o}}$, $C$ and $\delta$ by matching the initial value of the Stokes vector $\mathbf{S}\left(0\right)=\left(S_{1}\left(0\right),S_{2}\left(0\right),S_{3}\left(0\right)\right)$ with equations (32)-(34) evaluated at $s=0$.
In Fig. 8 we plot the above solutions for the following cases, assuming $r_{\mathrm{o}}=2r_{*}\approx 2\times 10^{6}\,\mathrm{cm}$, $\theta_{\mathrm{o}}=60^{\circ}$, $\alpha=45^{\circ}$, $\eta\approx 3.97\times 10^{-32}$ for $\hat{\Omega}_{1,2\left(\mathrm{min}\right)}$ (evaluated through (18), (19)) and $\omega\approx 2\pi\times 10^{18}\,\mathrm{Hz}$ ($k\approx 2.0958\times 10^{8}\,\mathrm{cm^{-1}}$) for the X-ray emissions, with the initial Stokes vector $\mathbf{S}\left(0\right)=\left(0.8,0,0.6\right)$:
@@
$\hat{\Omega}_{1\left(\mathrm{min}\right)}\approx-2.5511\times 10^{-16}$, $\hat{\Omega}_{2\left(\mathrm{min}\right)}\approx-2.2054\times 10^{-16}$, $s_{\dagger 1}\approx 1.0202\times 10^{6}\,\mathrm{cm}$, $s_{\dagger 2}\approx 7.5467\times 10^{5}\,\mathrm{cm}$, $p=100$, $\chi=0.999$, $S_{\mathrm{o}}\approx 0.7960$, $C\approx 0.6052$ and $\delta\approx 0.8537$,
@@
$\hat{\Omega}_{1\left(\mathrm{min}\right)}\approx-8.0001\times 10^{-15}$, $\hat{\Omega}_{2\left(\mathrm{min}\right)}\approx-3.9240\times 10^{-16}$, $s_{\dagger 1}\approx 1.0202\times 10^{6}\,\mathrm{cm}$, $s_{\dagger 2}\approx 7.5467\times 10^{5}\,\mathrm{cm}$, $p=100$, $\chi=0.999$, $S_{\mathrm{o}}\approx 0.6013$, $C\approx 0.7990$ and $\delta\approx 1.5056$,
@@
$\hat{\Omega}_{1\left(\mathrm{min}\right)}\approx-6.3776\times 10^{-13}$, $\hat{\Omega}_{2\left(\mathrm{min}\right)}\approx-2.7567\times 10^{-14}$, $s_{\dagger 1}\approx 1.0202\times 10^{6}\,\mathrm{cm}$, $s_{\dagger 2}\approx 7.5467\times 10^{5}\,\mathrm{cm}$, $p=100$, $\chi=0.056$, $S_{\mathrm{o}}\approx 0.6010$, $C\approx 0.7993$ and $\delta\approx 1.5133$.
These plots compare with Figs. 3b (for Example (i)), 4b (for Example (ii)) and 5b (for Example (iii)) in Section 2.2.1, respectively.
@@
Figure 8: Plots of the analytical solutions $S_{1}\left(s\right)$, $S_{2}\left(s\right)$ and $S_{3}\left(s\right)$, given the initial Stokes vector $\mathbf{S}\left(0\right)=\left(0.8,0,0.6\right)$.
@@
$\Psi\left(s;p=100\right)=k\sqrt{a_{1}^{2}+a_{2}^{2}}b^{-\frac{401}{200}}s^{\frac{1}{200}}e^{-\frac{1}{2}bs}\left[M_{\frac{1}{200},\frac{101}{200}}\left(bs\right)-\left(bs\right)^{\frac{201}{200}}e^{-\frac{1}{2}bs}\right]\approx 0.24k\sqrt{\hat{\Omega}_{1\left(\mathrm{min}\right)}^{2}+\hat{\Omega}_{2\left(\mathrm{min}\right)}^{2}}\pi s.$ (36)
Using this, we can estimate how much the oscillations for the three cases have progressed, for example, during $0\leq s\lesssim 1.2\times 10^{6}\,\mathrm{cm}$:
$\Psi\left(s\approx 1.2\times 10^{6}\,\mathrm{cm};p=100\right)$$\approx\left\{\begin{array}[]{ll}0.02\pi\hskip 9.0pt\text{(a fraction of an oscillation)}&\hskip 12.0pt\text{for Case (a)},\\ 0.5\pi\hskip 13.0pt\text{(about half an oscillation)}&\hskip 12.0pt\text{for Case (b)},\\ 39\pi\hskip 15.0pt\text{(multiple oscillations)}&\hskip 12.0pt\text{for Case (c)},\end{array}\right.$ (40)
which can be checked by comparison with Figs. 8a, 8b and 8c, respectively.
Recalling Example (i) from Section 2.2.1, one finds that the condition for perturbation can be equivalently expressed using (36):

### v3

@@ 2.2.2 Approximate analytical solutions
Plotting the birefringent functions $\hat{\Omega}_{1}\left(s\right)$ and $\hat{\Omega}_{2}\left(s\right)$, as given by (18) and (19), respectively, one can observe that they feature distinctive patterns; they can be well approximated by some analytic models, whose curves resemble the original plots. In Fig. 7 are plotted the birefringent functions for the three cases: (a) $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$), (b) $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$), (c) $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$) with solid lines (see Figs. 7a, 7b and 7c, respectively), where they have been evaluated with the same initial condition as assumed in Sect. 2.2.1.
@@
where $a_{1}>0$, $a_{2}<0$, $b>0$ and $p>0$ are free parameters; with suitable values chosen for these, our model functions can give rise to solutions of Eqs. (23)-(25) that match fairly well the numerical results obtained in Sect. 2.2.1. Here we express
$a_{1,2}=-\hat{\Omega}_{1\left(\mathrm{min}\right),2\left(\mathrm{max}\right)}\left[\frac{e}{\left(1-q\right)s_{\dagger 1}+qs_{\dagger 2}}\right]^{\frac{p+1}{p}},~{}~{}b=\frac{p+1}{p\left[\left(1-q\right)s_{\dagger 1}+qs_{\dagger 2}\right]},$ (31)
where $\hat{\Omega}_{1\left(\mathrm{min}\right),2\left(\mathrm{max}\right)}=\hat{\Omega}_{1,2}\left(s_{\dagger 1,2}\right)$, evaluated from (18), (19), with $s_{\dagger 1,2}$ denoting the extremum. We have set $p=100$ (a sufficiently large number) for all three cases, and $q=0.99999$ for (a) and (b), and $q=0.0532$ for (c) in Fig. 7.111111In particular, the values for $q$ have been chosen such that our solutions converge to the asymptotic limits that match well the numerical results given by Figs. 3b, 4b and 5b in Sect. 2.2.1, as $s$ tends to $\infty$.
@@
Figure 7: Plots of the birefringent functions $\hat{\Omega}_{1}\left(s\right)$ and $\hat{\Omega}_{2}\left(s\right)$ and their approximate analytic models $\hat{\Omega}_{1\left(\mathrm{m}\right)}\left(s\right)$ and $\hat{\Omega}_{2\left(\mathrm{m}\right)}\left(s\right)$.
Plugging Eq. (30) into Eqs. (23)-(25), we obtain analytical solutions as follows (for a complete derivation, see Appendix B):
@@
and $M_{\kappa,\mu}\left(z\right)$ denotes a Whittaker function of the first kind, [^12] The expression inside the square brackets in Eq. (35) has been reduced from its original form as given by Eq. (83) in Appendix B, using the identity $M_{\left(2p+1\right)/\left(2p\right),\left(p+1\right)/\left(2p\right)}\left(bs\right)=\left(bs\right)^{\left(2p+1\right)/\left(2p\right)}e^{-bs/2}M\left(0,\left(2p+1\right)/p,bs\right)$, with the Kummer function $M\left(0,\left(2p+1\right)/p,bs\right)=1$ as a special case NIST2024 . and $a_{1,2}$ and $b$ are given by (31). Here we determine $S_{\mathrm{o}}$, $C$ and $\delta$ by matching the initial value of the Stokes vector $\mathbf{S}\left(0\right)=\left(S_{1}\left(0\right),S_{2}\left(0\right),S_{3}\left(0\right)\right)$ with Eqs. (32)-(34) evaluated at $s=0$.
In Fig. 8 we plot the above solutions (32)-(34) for the following cases, assuming $r_{\mathrm{o}}=2r_{*}\approx 2\times 10^{6}\,\mathrm{cm}$, $\theta_{\mathrm{o}}=60^{\circ}$, $\alpha=45^{\circ}$, $\eta\approx 3.97\times 10^{-32}$ for $\hat{\Omega}_{1\left(\mathrm{min}\right),2\left(\mathrm{max}\right)}=\hat{\Omega}_{1,2}\left(s_{\dagger 1,2}\right)$ (evaluated via (18) and (19)) and $\omega\approx 2\pi\times 10^{18}\,\mathrm{Hz}$ ($k\approx 2.0958\times 10^{8}\,\mathrm{cm^{-1}}$) for the X-ray emissions, with the initial Stokes vector $\mathbf{S}\left(0\right)=\left(0.8,0,0.6\right)$:
@@
$\hat{\Omega}_{1\left(\mathrm{min}\right)}\approx-2.5511\times 10^{-16}$, $\hat{\Omega}_{2\left(\mathrm{max}\right)}\approx 2.2054\times 10^{-16}$, $s_{\dagger 1}\approx 1.0202\times 10^{6}\,\mathrm{cm}$, $s_{\dagger 2}\approx 7.5467\times 10^{5}\,\mathrm{cm}$, $p=100$, $q=0.99999$ for $a_{1,2}$ and $b$ (evaluated via (31)), and $S_{\mathrm{o}}\approx 0.7961$, $C\approx 0.6052$, $\delta\approx 2.2879$,
@@
$\hat{\Omega}_{1\left(\mathrm{min}\right)}\approx-8.0001\times 10^{-15}$, $\hat{\Omega}_{2\left(\mathrm{max}\right)}\approx 3.9240\times 10^{-16}$, $s_{\dagger 1}\approx 1.0202\times 10^{6}\,\mathrm{cm}$, $s_{\dagger 2}\approx 7.5467\times 10^{5}\,\mathrm{cm}$, $p=100$, $q=0.99999$ for $a_{1,2}$ and $b$ (evaluated via (31)), and $S_{\mathrm{o}}\approx 0.6013$, $C\approx 0.7990$, $\delta\approx 1.6360$,
@@
$\hat{\Omega}_{1\left(\mathrm{min}\right)}\approx-6.3776\times 10^{-13}$, $\hat{\Omega}_{2\left(\mathrm{max}\right)}\approx 2.7567\times 10^{-14}$, $s_{\dagger 1}\approx 1.0202\times 10^{6}\,\mathrm{cm}$, $s_{\dagger 2}\approx 7.5467\times 10^{5}\,\mathrm{cm}$, $p=100$, $q=0.0532$ for $a_{1,2}$ and $b$ (evaluated via (31)), and $S_{\mathrm{o}}\approx 0.6010$, $C\approx 0.7993$, $\delta\approx 1.6283$.
These plots compare with Figs. 3b (for Example (i)), 4b (for Example (ii)) and 5b (for Example (iii)) in Sect. 2.2.1, respectively.
@@
Figure 8: Plots of the analytical solutions $\mathbf{S}_{\left(\mathrm{a}\right)}\left(s\right)=\left(S_{1\left(\mathrm{a}\right)}\left(s\right),S_{2\left(\mathrm{a}\right)}\left(s\right),S_{3\left(\mathrm{a}\right)}\left(s\right)\right)$, given the initial Stokes vector $\mathbf{S}\left(0\right)=\left(0.8,0,0.6\right)$.
@@
$\Psi\left(s;p=100\right)$ $=k\sqrt{a_{1}^{2}+a_{2}^{2}}b^{-\frac{401}{200}}s^{\frac{1}{200}}e^{-\frac{1}{2}bs}\left[M_{\frac{1}{200},\frac{101}{200}}\left(bs\right)-\left(bs\right)^{\frac{201}{200}}e^{-\frac{1}{2}bs}\right]$
$\approx 0.24k\sqrt{\hat{\Omega}_{1\left(\mathrm{min}\right)}^{2}+\hat{\Omega}_{2\left(\mathrm{max}\right)}^{2}}\pi s.$ (36)
Using this, we can estimate how much the oscillations for the three cases have progressed, for example, during $0\leq s\lesssim 2.5\times 10^{6}\,\mathrm{cm}$:
$\Psi\left(s\approx 2.5\times 10^{6}\,\mathrm{cm};p=100\right)\approx\left\{\begin{array}[]{ll}0.04\pi\hskip 3.0pt\text{(a fraction of an oscillation)}&\hskip 3.0pt\text{for Case (a)},\\ \pi\hskip 8.0pt\text{(about half an oscillation)}&\hskip 3.0pt\text{for Case (b)},\\ 80\pi\hskip 11.0pt\text{(multiple oscillations)}&\hskip 3.0pt\text{for Case (c)},\end{array}\right.$ (40)
each of which can be checked by comparison with Figs. 8a, 8b and 8c, respectively.
Recalling Example (i) from Sect. 2.2.1, one finds that the condition for perturbation can be equivalently expressed with the help of (36):
$\left|k\hat{\Omega}_{1,2}\left(s\right)s\right|_{\mathrm{max}}\sim\left|k\hat{\Omega}_{1,2}\left(s_{\dagger 1,2}\right)s_{\dagger 1,2}\right|$ $\sim 0.24k\sqrt{\hat{\Omega}_{1\left(\mathrm{min}\right)}^{2}+\hat{\Omega}_{2\left(\mathrm{max}\right)}^{2}}\pi\times 10^{6}\,\mathrm{cm}$